ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 34B Sadyrbaev, Felix, Institute of Mathematics and Computer Science, Riga, Latvia Nonlinear Sturm type theorems Two nontrivial solutions of a second order nonlinear equation

x′′ + f(t, x, x′) = 0

are considered. Zero separation results are presented in terms of corresponding equations of variations. Classical Sturm theorem for linear equations is included.